OpenStudy (anonymous):

derivative f(x)= 3e^(-7x^2)

OpenStudy (anonymous):

chain rule!

jimthompson5910 (jim_thompson5910):

use the chain rule to get f(x) = 3*e^(-7x^2) f ' (x) = d/dx[ 3*e^(-7x^2) ] f ' (x) = 3 * d/dx[ e^(-7x^2) ] f ' (x) = 3e^(-7x^2) * d/dx[ (-7x^2) ] f ' (x) = 3e^(-7x^2) * (-14x) f ' (x) = -42x*e^(-7x^2)

OpenStudy (anonymous):

to find critical point f ' (x) = -42x*e^(-7x^2) =0 and it would be zero?

jimthompson5910 (jim_thompson5910):

yep because you use the zero product property to get -42x = 0 or e^(-7x^2) =0 but e^(-7x^2) =0 has no solutions

OpenStudy (anonymous):

Great! thanks. do you know what inflection point ? I have never heard in the class

jimthompson5910 (jim_thompson5910):

it's where the concavity flips

OpenStudy (anonymous):

from the same question that you helped me, would it be zero again?

jimthompson5910 (jim_thompson5910):

well you need to find f '' to find the intervals of concavity and the inflection point

OpenStudy (anonymous):

oh. I see.

OpenStudy (anonymous):

do i use product rule first?

OpenStudy (anonymous):

I got 41e^(-7x^2)(14x^2-1)

OpenStudy (anonymous):

-1/sqrt(14),1/sqrt(14)

OpenStudy (anonymous):

for the points?

jimthompson5910 (jim_thompson5910):

I think you meant 42 instead of 41

jimthompson5910 (jim_thompson5910):

I got 42*e^(-7x^2)*(14x^2-1)

jimthompson5910 (jim_thompson5910):

for f ''